NASA ISS On-Orbit Status 17 December 2013

©NASA

NASA ISS On-Orbit Status 17 December 2013

Today: USOS EVA and Orbital Status: A Program decision was made to pursue USOS EVAs on December 21st, 23rd, and 25th (if required) in order to repair the ETCS Loop A Pump Module. The EVAs will be carried out using EMUs 3010 and 3011. Orbital launch will be delayed until after the beta cutout in January.

External Thermal Control System (ETCS) Loop A Status: Overnight ground teams performed a Lab Low Temperature Loop (LTL) Interface Heat Exchanger (IFHX) integration test. During this test the team used new tools developed by the Flight Control Team that allow improved control of Radiator Return Isolation Valve movement. More precise control of this valve provides enhanced regulation of the ETCS Loop A temperatures. During the test, the isolated ETCS was successfully re-integrated with the Lab LTL through the IFHX. The team then began to increase the heat load on the internal side of the LTL IFHX while attempting to balance the added heat load on ETCS loop A by manipulating the Radiator Return Isolation valve. After approximately 40 minutes the pump shut down when the freeze protection Fault Detection, Isolation, and Recovery (FDIR) limit was reached. Subsequent attempts to stabilize the Loop A temperature and reintegrate the loop were unsuccessful.

Ground teams determined that the current method of commanding the Radiator Return Isolation valve does not provide sufficient control, and additional precision is required to maintain stable temperatures. A software patch to the S1-1 MDM was delivered and loaded this afternoon which should allow for finer control over the Radiator Return Isolation Valve positioning. Additional external loop temperature control tests will be performed this evening followed by reintegration attempts. Due to the loss of heat rejection for the Node 2 Moderate Temperature Loop (MTL), half of the loads in the Node 2, JEM and Columbus modules remain powered off.

Water Process Assembly (WPA) Status: Overnight ground teams performed WPA troubleshooting which resulted in the MLS successfully starting up without an overspeed fault. Following this test, the crew reconnected process line B between the two Water Recovery System (WRS) racks. Currently the WPA is in standby and heating up in preparation for future processing. The first WPA process run will likely not occur until tomorrow at the earliest.

OGA Hydrogen Sensor ORU Removal and Replacement (R&R): Today the crew completed an R&R of the Oxygen Generation System (OGS) Hydrogen Sensor ORU. This task was being done as a preventative maintenance item that is performed approximately every 150 days. During the activity the crew reported that foam ring installed around the OGA O2 outlet port was crumbling and falling apart. The purpose of this foam ring is to cushion the port against the rack door to ensure that O2 is expelled into the ISS cabin and prevent O2 from building up behind the rack. Currently, the OGA is powered down and the H2 sensor Quick Disconnects (QD) are not connected while the ground team reviews photos of the degraded foam ring and determines the appropriate forward path. There is currently one spare O2 outlet ORU onboard.

Extravehicular Activity (EVA) Preparations: In preparation for a potential contingency EVA, Flight Engineer (FE)-3 Hopkins and FE-5 Mastracchio performed a variety of EVA related preparation activities. Today's activities included a review of the Hard Upper Torso (HUT)/Secondary Oxygen Packs (SOP) swap procedures, QD training, EVA procedures and training material review.

On-board Training (OBT) Robotics Onboard Trainer (ROBoT) Session 1: Utilizing the ROBOT Training Simulator, FE-6 Wakata gained proficiency in Space Station Remote Manipulator System (SSRMS) Operations associated with an ETCS Loop A Pump Module removal and replacement.

Today's Planned Tasks

Combustion Integrated Rack Alignment Guide Removal [Complete]
EMU Long Life Battery (LLB) Charge Terminate [Complete]
Replacement of Urine Receptacle (МП) and Filter-Insert (Ф-В) in АСУ Install [Complete]
Locating EVA hardware and tools. [Complete]
Extravehicular Activity (EVA) Procedure Review [Complete]
Hand ergometry. [Complete]
Dynamic Onboard Ubiquitous Graphics (DOUG) Software Review [Complete]
СОЖ Maintenance [Complete]
MOTOCARD. Experiment Ops [Complete]
IDENTIFIKATSIYA. Copy ИМУ-Ц micro-accelerometer data to laptop [Complete]
Comm assets configuration for conversations from Soyuz #710 [Complete]
Troubleshooting Noise on ТПК 710 ИнПУ2 in Display+TV mode [Complete]
Crew sets up an HD view from the Lab AVN 443 Encoder[Complete]
Microbial Removal Filter Flow [Complete]
Equipment Lock (E-LK) Preparation [Complete]
Recharging КОБ2 Loop with Coolant from DC1 Hydraulic Loop Accumulator [Complete]
EMU On-orbit Fitcheck Verification [Complete]
Robotics Extravehicular Activity (EVA) Procedure Conference [Complete]
EMU Wet Helmet Absorption Pad (HAP) Familiarization [Complete]
Anti-virus base upgrade on laptops ВКС [Complete]
Comm back to nominal [Complete]
СТТС comm system switchover to the primary set [Complete]
Locating EVA hardware and tools [Complete]
IMS Update[Complete]
ISS repress using air from section 1 of ТКГ 420 СрПК [Complete]
ISS repress using air from ТКГ 420 СрПК Section 1 [Complete]
Task List

ARED Cylinder Evacuation
OGA Hydrogen ORU Purge Adapter (HOPA) Operations [Complete]
ТКГ 421 Transfers and IMS Updates /Transfer Ops
ЕДВ-У R&R
Preparation of reports for Roskosmos site
URAGAN. Observations and Photography
ECON-M. Observations and Photography
OGA Hydrogen ORU Purge Adapter (HOPA) Operations, Hydrogen (H2) Sensor ORU R&R, and Air Avionics Assembly (AAA) Cleaning
National Geographic B-Role
EVA Camera Configuration
PAO Event
COL D1 FD2 MSR
Cygnus Power Jumper Build [Complete]
DOS3D PDP Check
IFM stowage Locate
Kubik 6 Belt Search
National Engr Week Message
P/TV Trash
PPS Wipe
Pre-gather COL A/L
Pre-gather Empty CTBs
SM PCS HD Swap Imaging
WINEARTH HISD Photos
Ground

Nominal Commanding [Complete]
EXT MDM Swap [Complete]
S1-1 MDM Software Patch Load [Complete]
Three-Day Look Ahead:

Tuesday, 12/17: EVA Tool Gather, MDCA Hardware Replace, Body Measurement, Tonometry
Wednesday, 12/18: IMV Flow measurements, OCT, SPDA) Jumper Assembly and Installation
Thursday, 12/19: TBD
QUICK ISS Status - Environmental Control Group:

Component - Status
Elektron - On
Vozdukh - Manual
[СКВ] 1 - SM Air Conditioner System ("SKV1") - On
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Off
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) - Stop
Urine Processing Assembly (UPA) - Standby
Trace Contaminant Control System (TCCS) Lab - Off
Trace Contaminant Control System (TCCS) Node 3 - Full up

Please follow SpaceRef on Twitter and Like us on Facebook.


  • submit to reddit